# Diffusion Models in 3D Vision: A Survey

### v1

@@ IV-C Text to 3D
Nichol et al.[72] is one of the first to explore how to use DDPMs to efficiently generate 3D point clouds from natural language descriptions. The authors first use a text-to-image diffusion model to generate a single synthetic view, followed by a second conditional diffusion model to produce the 3D point cloud based on the generated image.
Li et al.[73] introduce a novel 3D modeling framework called Diffusion-SDF, aimed at addressing the challenging task of text-to-shape synthesis. The authors propose a patch-based SDF autoencoder that maps the truncated signed distance fields (TSDFs) of 3D shapes into local Gaussian latent representations. And they introduce a voxelized diffusion model to capture local information and the relationships between patches (patch-to-patch) and the global structure (patch-to-global).
Metzer et al.[74] present a novel NeRF model called Latent-NeRF, which operates in latent space instead of the standard RGB space. This approach eliminates the burden of encoding the rendered RGB images into latent space at each guidance step. The authors utilize text features as another input in the latent space for text-based guidance, enhancing the model’s ability to generate 3D representations based on textual descriptions.
Chen et al.[75] introduce a method called Fantasia3D, which achieves high-quality conversion from text prompts to 3D content creation by decoupling the modeling of geometry and appearance. Text features are specifically employed in the modeling of both geometry and appearance, guiding the generation process to produce the desired outcomes effectively.
Wu et al.[76] introduce a model named STPD, which generates 3D shapes and colors through a staged diffusion process. Initially, features from text and sketches are extracted separately. Multi-head attention is then employed to fuse the features from the sketch and text. Following this, a staged diffusion model is used to first generate the shape, and subsequently, colors are assigned based on the shape generated in the first stage.
Poole et al.[77] explore how to utilize a pretrained 2D text-to-image diffusion model to synthesize 3D models without requiring 3D training data. They employ a pretrained 2D text-to-image diffusion model as a prior and propose a loss function based on probabilistic density distillation. This approach allows the use of the 2D diffusion model as an prior. Finally, they optimize a randomly initialized 3D NeRF through gradient descent, aiming to achieve low loss from 2D images rendered from random viewpoints. This approach effectively addresses the challenges of scarce 3D data and the inefficiencies of 3D denoising architectures, presenting a novel possibility for 3D content creation.
Lu et al.[78] propose a novel multi-view 2.5D diffusion model, which fine-tunes a pretrained 2D model to directly capture the structural distribution of 3D data while retaining the strong generalization capabilities of the original 2D diffusion model. This approach addresses the limitations of current methods, which either rely on pretrained 2D diffusion models with time-consuming SDS or directly use 3D diffusion models constrained by limited 3D data, sacrificing diversity in generation.
@@
Seo et al.[79] investigate the issue that 2D diffusion models lack an understanding of 3D space, often resulting in distorted and incoherent 3D scenes. To tackle this, the authors propose a new framework called 3DFuse. The framework first generates a rough 3D point cloud based on text prompts and projects it into arbitrary views to create rough depth maps. These depth maps serve as explicit 3D consistency conditions, which are fed into the 2D diffusion model. This enables the model to optimize each viewpoint in a 3D-aware manner, improving the coherence and structure of the generated 3D scenes.
Wang et al.[80] propose a novel method called Variational Score Distillation (VSD). VSD is a particle-based variational framework that models 3D parameters as random variables, contrasting with the constants used in traditional Score Distillation Sampling (SDS). By optimizing the distribution of the 3D scene, VSD aims to make the image distribution rendered from all views closely match the distribution defined by the pretrained 2D diffusion model. This approach enhances the flexibility and accuracy of 3D content generation by leveraging the strengths of both variational inference and diffusion models.
Bahmani et al.[81] introduce a technique called Hybrid Score Distillation Sampling (SDS). This method employs an alternating optimization process, beginning with the generation of an initial static 3D scene using a 3D-aware text-to-image model that avoids the “Janus problem”. Following this, the process alternates between variational SDS and text-to-image model supervision to enhance the scene’s appearance. Finally, the technique incorporates video SDS and text-to-video model supervision to introduce motion into the scene, resulting in a dynamic and visually compelling 3D representation.
Huang et al.[82] identify a conflict between the 3D optimization process and the uniform timestep sampling in diffusion models, which led to slow convergence, poor quality of generated 3D models, and low diversity in generation. To address these issues, the paper proposes a Time Prioritized Score Distillation Sampling (TP-SDS) strategy. This approach uses a monotonically decreasing function to prioritize different diffusion timesteps during the 3D optimization process, aligns the 3D optimization process more effectively with the sampling steps.
Chen et al.[83] introduce GSGEN, a novel method for text-to-3D generation using 3D Gaussian scattering. The method employs a two-stage optimization strategy. In the first stage, it utilizes joint guidance from 2D and 3D diffusion priors to shape a 3D-consistent coarse structure. In the second stage, during appearance refinement, the model enhances details through a compactness-based densification process
@@
Ling et al.[84] propose a new method for generating high-quality 4D dynamic scenes from textual descriptions by leveraging dynamic 3DGS and composite diffusion models. The authors use dynamic 3D Gaussian Mixtures and deformation fields as the backbone for 4D representation. The deformation fields capture scene dynamics and transform the set of 3D Gaussians to represent object motion. The approach combines text-to-image, text-to-video, and 3D-aware multi-view diffusion models, providing feedback within a score distillation-based synthesis framework.
Yi et al.[85] explore how to efficiently generate 3D Gaussian representations from text prompts while maintaining 3D consistency and detailed generation. The authors propose a framework called GaussianDreamer, which bridges 3D and 2D diffusion models through Gaussian splatting. In this framework, the 3D diffusion model provides an initialization prior, and the 2D diffusion model enhances both the geometry and appearance, resulting in a detailed and coherent 3D output.
@@ IV-D Texture Generation
Mesh generation involves creating polygonal meshes that define the structure of 3D objects, while texture generation refers to mapping realistic surface details onto those meshes. Diffusion models aid in both processes by generating high-fidelity meshes and textures from learned distributions. These models can ensure geometric accuracy and surface detail that align with real-world appearances, benefiting applications in graphics, simulation, and AR/VR environments.
Cao et al.[86] introduce TexFusion, which achieves text-guided 3D texture synthesis by using a latent diffusion model and a novel 3D-consistent diffusion model sampling scheme. This method runs a 2D denoising network across different 2D rendered views and aggregates the predicted results into a latent texture map after each denoising iteration. This approach effectively generates textures with rich details and consistency, responding to text prompts, offering new possibilities for AI-based texture generation for 3D assets.
Chen et al.[87] employ a pretrained depth-aware image diffusion model, which can generate high-resolution partial textures based on text prompts. Additionally, the authors progressively synthesize high-resolution local textures from different viewpoints to avoid the accumulation of inconsistencies and stretched artifacts between views.
Yu et al.[88] investigate methods for automatically generating high-fidelity and geometry-compatible texture images on 3D mesh models. The authors propose Point-UV Diffusion, a two-stage coarse-to-fine framework that combines denoising diffusion models with UV mapping. First, a point diffusion model is used to synthesize low-frequency texture components, introducing a style-guided mechanism to address color distribution bias in the dataset. These colored points are then mapped to 2D UV space, and 3D coordinate interpolation is used to generate a rough texture image, which maintains 3D consistency and continuity.
Gupta et al.[89] propose an architecture called 3DGen, which combines a VAE and a conditional diffusion model to generate high-quality textured and non-textured 3D meshes. It consists of two stages: first, a Triplane VAE encodes the point cloud from the input mesh into a triplane Gaussian latent space, learning to reconstruct continuous textured meshes from this latent space. Next, a diffusion model is trained to generate triplane features.
Richardson et al.[90] use a pretrained depth-to-image diffusion model to render 3D models from different viewpoints, incorporating a dynamic trimap segmentation technique. By considering trimap segmentation, it proposes an improved diffusion process for generating seamless textures across various views. This enables text-guided texture generation, transfer, and editing for 3D models, significantly enhancing the quality and speed of texture generation.
Chen et al.[91] introduce SceneTex, which aims to address the problem of texture synthesis for indoor scenes. Unlike previous methods, SceneTex formulates the texture synthesis task as an optimization problem in RGB space. By employing a multi-resolution texture field and a cross-attention decoder, it effectively leverages the depth-to-image diffusion prior to generate high-quality, style-consistent textures for indoor scenes.
Li et al.[92] aim solve the infinite possibilities of UV mapping problem when generating 2D texture maps in traditional 2D diffusion methods, while using point clouds helps to avoid variations caused by different mesh topologies and UV mappings. They proposes an octree-based diffusion model (TexOct), which uses an octree structure to efficiently represent the point cloud sampled from the surface of a given 3D model. This method directly optimizes texture generation in 3D space, effectively addressing texture errors caused by self-occlusion and texture degradation due to sparse sampling.
Yeh et al.[93] research how to transfer highly detailed and complex textures from real-world environments onto arbitrary objects. They introduce TextureDreamer, a novel framework by combining personalized modeling, score distillation, and geometry guidance, enabling it to extract texture information from a small number of images and transfer it onto target objects with different geometries.
Zeng et al.[94] introduce Paint3D, which aims to address the challenge of generating high-quality textures without embedded lighting information. By utilizing a pretrained 2D image diffusion model and a specially designed UV completion and upscaling diffusion model, Paint3D achieves the goal of generating high-quality, lighting-free 3D texture maps from text or image inputs.
Wang et al.[95] propose a novel convolution and DDPM-based reconstruction model, CRM, which leverages geometric priors and an efficient network architecture to achieve fast generation of high-fidelity 3D textured meshes from a single image.
@@
Zhang et al.[96] propose a novel approach to achieve multi-view consistency. This method is based on a pre-trained diffusion model, such as DDIM, which decodes the predicted noise-free latent states into color space at each denoising step and blends them to form a texture image. An optimization process is then used to update the latent codes, ensuring that the images rendered from the blended texture image remain consistent across all views.
@@
Zhang et al.[97] propose a new diffusion model that takes into account the given lighting environment during training to generate shading results under specific lighting conditions. By applying the same ambient lighting during the material extraction process, DreamMat can produce high-quality PBR materials that are consistent with the given geometry and free from any baked shadow effects.
@@ IV-E Human Avatar Generation
Wang et al.[98] investigate how to use diffusion models to generate high-fidelity 3D digital avatars, represented in NeRFs. Their RODIN model effectively generates detailed 3D digital avatars by unfolding 3D feature maps onto a 2D plane and applying 3D-aware convolutions which accounts for the projection relationship of 3D data onto the 2D plane.
Jiang et al.[99] propose a method called AvatarCraft, which uses diffusion models as guidance to generate 3D human avatars with parameterized shape and pose control based on text prompts. By mapping the target character mesh to a template character mesh through an explicit deformation field, the neural implicit field is deformed, simplifying the animation and deformation of the generated avatars. This method also supports high-quality novel view synthesis, scene composition, and animation.
Chen et al.[100] present a novel 3D deformation diffusion model capable of generating 3D mesh models based on input parameters such as identity and expression. By combining the 3D deformation model with a multi-view consistency diffusion approach, the model achieves the generation of controllable and realistic 3D avatars from a single image.
Lei et al.[101] propose the DiffusionGAN3D framework, which combines a pretrained 3D generative model (such as EG3D) with a text-to-image diffusion model. The framework employs Relative Distance Loss to enhance diversity during the domain adaptation process and utilizes case-specific learnable triplanes to improve the network’s generative capabilities.
Li et al.[92] address several key issues present in existing facial texture generation methods, including low texture quality, loss of identity information, and inadequate handling of occlusions. To tackle these challenges, the authors propose UV-IDM, which leverages the powerful texture generation capabilities of latent diffusion models (LDM) to generate realistic facial textures based on the Basel Face Model (BFM). To preserve identity information during the reconstruction process, an Identity Condition Module (ICM) is designed, allowing any outdoor image to serve as a robust condition for the LDM to guide texture generation.
Zhou et al.[102] introduce a new method called UltrAvatar, which uses Diffuse Color Extraction Model and Authenticity Guided Texture Diffusion Model to generate realistic and animatable 3D avatars from text prompts or a single image.
Chen et al.[103] introduce the first diffusion model-based framework for 3D human generation called PrimDiffusion. The authors propose a 3D human representation method based on volumetric primitives, modeling the human body as multiple small volumes with radiance and motion information. By directly manipulating the denoising and diffusion processes on the volumetric primitives, the framework addresses the high computational cost of 3D representations and the complexity of 3D human topology.
Kim et al.[104] present a 3D human generation pipeline called Chupa, aimed at addressing the challenge of generating realistic 3D human meshes with diverse identities, poses, and random details. The researchers break the problem down into 2D normal map generation and normal map-based 3D reconstruction, improving computational efficiency and generating multi-view consistent 3D models. First, a pose-conditioned diffusion model is used to simultaneously generate realistic front and back normal maps of the character. Then, mesh optimization is employed to “sculpt” an initial SMPL-X mesh to match the details of the normal maps
Svitov et al.[105] propose a method called DINAR, which creates realistic and controllable 3D cartoon characters from a single image while maintaining rendering quality and good generalization to new poses and viewpoints. DINAR combines neural textures with the SMPL-X human model, achieving realistic quality while keeping the cartoon characters easy to manipulate and enabling fast inference. A latent diffusion model is used to recover the textures.
Kolotouros et al.[106] propose a method for generating realistic and animatable 3D human avatars solely from text descriptions. A statistical human model is then incorporated to enhance the realism of the human body structure, appearance, and deformations. Additionally, the quality of the generated models is improved by adding details to perceptually important body areas, such as the face and hands.
Ho et al.[107] introduces a new 3D human reconstruction pipeline called SiTH, which addresses the challenge of reconstructing realistic and detail-rich 3D human models from a single-view image. The SiTH method breaks down the single-view reconstruction problem into two subproblems: generating the illusion of a back view image and mesh reconstruction. First, an image-conditioned diffusion model is used to predict the appearance of the back view of the input image. Then, using both the foreground and back view images as guidance, the method reconstructs a textured mesh of the entire body.
Huang et al.[108] present a new method called HumanNorm, which enhances the model’s 2D perception of 3D geometry by learning a normal-adapted diffusion model and a normal-aligned diffusion model. Additionally, it introduces a progressive geometry generation strategy and a multi-step Score Distillation Sampling (SDS) loss function to improve the performance of 3D human generation.
Stathopoulos et al.[109] present a method called ScoreHMR, aimed at addressing the problem of predicting human model parameters from 2D image evidence. The ScoreHMR method mimics model fitting procedure, where a diffusion model is trained to capture the conditional distribution of human model parameters for a given input image. By using task-specific scores to guide its denoising process, ScoreHMR effectively resolves inverse problems across various applications without the need to retrain the diffusion model for each specific task.
Xu et al.[110] introduce a new task of predicting human-object interactions (HOI) in 3D space, with the primary goal of generating long-term, diverse, and vivid 3D HOI sequences. The authors propose a framework called InterDiff, which combines interaction diffusion with physics-informed interaction correction to generate realistic and physically plausible future 3D sequences of human-object interactions.
@@
Yuan et al.[111] introduce a novel physics-guided human motion diffusion model called PhysDiff, which addresses the issue that traditional motion diffusion models often neglect physical laws, leading to the generation of physically unrealistic motions. The model incorporates physical constraints during the diffusion process by using a physics-based motion projection module. This module employs motion imitation from a physics simulator to project the denoised motions in the diffusion steps onto physically plausible movements.

### v2

@@ IV-C Text to 3D
Among all the researches, Nichol et al.[71] is one of the first to explore how to use DDPMs to efficiently generate 3D point clouds from natural language descriptions. The authors first use a text-to-image diffusion model to generate a single synthetic view, followed by a second conditional diffusion model to produce the 3D point cloud based on the generated image. Subsequently, Li et al.[72] propose a patch-based SDF autoencoder that maps the truncated signed distance fields (TSDFs) of 3D shapes into local Gaussian latent representations. And they introduce a voxelized diffusion model to capture local information and the relationships between patches (patch-to-patch) and the global structure (patch-to-global). To eliminates the burden of encoding the rendered RGB images into latent space at each guidance step, Metzer et al.[73] present Latent-NeRF, which operates in latent space instead of the standard RGB space. Moreover, Seo et al.[74] investigate the issue that 2D diffusion models lack an understanding of 3D space, often resulting in distorted and incoherent 3D scenes. They generates a rough 3D point cloud based on text prompts and projects it into arbitrary views to create rough depth maps, serving as explicit 3D consistency conditions and are fed into the 2D diffusion model. This enables the model to optimize each viewpoint in a 3D-aware manner, improving the coherence and structure of the generated 3D scenes.
To generate more accurate results, some studies employ a two-stage approach to generate results. Chen et al.[75] decouple the modeling of geometry and appearance where text features are specifically employed in the modeling of both geometry and appearance, guiding the generation process to produce the desired outcomes effectively. And Wu et al.[76] use a staged diffusion model in which they first generate the shape, and subsequently, and then colors are assigned based on the shape generated in the first stage.
@@
For 3DGS-based [39] generation, Chen et al.[77] introduce GSGEN, a novel method for text-to-3D generation using 3D Gaussian scattering. The method employs a two-stage optimization strategy. In the first stage, it utilizes joint guidance from 2D and 3D diffusion priors to shape a 3D-consistent coarse structure. In the second stage, during appearance refinement, the model enhances details through a compactness-based densification process. Ling et al.[78] propose a new method for generating high-quality 4D dynamic scenes from textual descriptions by leveraging dynamic 3DGS and composite diffusion models. The authors use dynamic 3D Gaussian Mixtures and deformation fields as the backbone for 4D representation. The deformation fields capture scene dynamics and transform the set of 3D Gaussians to represent object motion. The approach combines text-to-image, text-to-video, and 3D-aware multi-view diffusion models, providing feedback within a score distillation-based synthesis framework. And Yi et al.[79] explore how to efficiently generate 3D Gaussian representations from text prompts while maintaining 3D consistency and detailed generation. The authors propose a framework called GaussianDreamer, which bridges 3D and 2D diffusion models through Gaussian splatting. In this framework, the 3D diffusion model provides an initialization prior, and the 2D diffusion model enhances both the geometry and appearance, resulting in a detailed and coherent 3D output.
In addition to making improvements during the generation process, some methods attempt to enhance traditional Score Distillation Sampling (SDS) to improve the performance of generative models. Wang et al.[80] propose Variational Score Distillation (VSD), which is a particle-based variational framework that models 3D parameters as random variables, contrasting with the constants used in traditional SDS. By optimizing the distribution of the 3D scene, VSD aims to make the image distribution rendered from all views closely match the distribution defined by the pretrained 2D diffusion model. Bahmani et al.[81] introduce Hybrid Score Distillation Sampling (HSDS). This method employs an alternating optimization process, beginning with the generation of an initial static 3D scene using a 3D-aware text-to-image model that avoids the “Janus problem”. Following this, the process alternates between variational SDS and text-to-image model supervision to enhance the scene’s appearance. Furthermore, Huang et al.[82] identify a conflict between the 3D optimization process and the uniform timestep sampling in diffusion models, which led to slow convergence, poor quality of generated 3D models, and low diversity in generation. Thus, they proposes a Time Prioritized Score Distillation Sampling (TP-SDS) strategy, which uses a monotonically decreasing function to prioritize different diffusion timesteps during the 3D optimization process, aligns the 3D optimization process more effectively with the sampling steps.
To address the issue of low diversity in generated objects caused by the limited number of training samples for 3D tasks, some researchers have attempted to generate 3D objects using only 2D pretrained models without using any direct 3D samples. Poole et al.[83] explore how to utilize a pretrained 2D text-to-image diffusion model to synthesize 3D models without requiring 3D training data. They employ a pretrained 2D text-to-image diffusion model as a prior and propose a loss function based on probabilistic density distillation. This approach allows the use of the 2D diffusion model as an prior. Finally, they optimize a randomly initialized 3D NeRF through gradient descent, aiming to achieve low loss from 2D images rendered from random viewpoints. This approach effectively addresses the challenges of scarce 3D data and the inefficiencies of 3D denoising architectures, presenting a novel possibility for 3D content creation. Lu et al.[84] propose a novel multi-view 2.5D diffusion model, which fine-tunes a pretrained 2D model to directly capture the structural distribution of 3D data while retaining the strong generalization capabilities of the original 2D diffusion model. This approach addresses the limitations of current methods, which either rely on pretrained 2D diffusion models with time-consuming SDS or directly use 3D diffusion models constrained by limited 3D data, sacrificing diversity in generation.
@@ IV-D Texture Generation
Texture generation refers to mapping realistic surface details onto those meshes. Diffusion models aid in both processes by generating high-fidelity textures from learned distributions. These models can ensure geometric accuracy and surface detail that align with real-world appearances, benefiting applications in graphics, simulation, and AR/VR environments.
Cao et al.[85] introduce TexFusion, which achieves text-guided 3D texture synthesis by using a latent diffusion model and a novel 3D-consistent diffusion model sampling scheme. This method runs a 2D denoising network across different 2D rendered views and aggregates the predicted results into a latent texture map after each denoising iteration. This approach effectively generates textures with rich details and consistency, responding to text prompts, offering new possibilities for AI-based texture generation for 3D assets. Yu et al.[86] investigate methods for automatically generating high-fidelity and geometry-compatible texture images on 3D mesh models. The authors propose Point-UV Diffusion, a two-stage coarse-to-fine framework that combines denoising diffusion models with UV mapping. Gupta et al.[87] combine a VAE and a conditional diffusion model to generate high-quality textured 3D meshes. They use a Triplane VAE to encode the point cloud from the input mesh into a triplane Gaussian latent space and then learn to reconstruct continuous textured meshes from this latent space. Li et al.[88] propose an octree-based diffusion model (TexOct), which uses an octree structure to efficiently represent the point cloud sampled from the surface of a given 3D model. This method directly optimizes texture generation in 3D space, effectively addressing texture errors caused by self-occlusion and texture degradation due to sparse sampling. Yeh et al.[89] combine personalized modeling, score distillation, and geometry guidance to extract texture information from a small number of images and transfer it onto target objects with different geometries. Wang et al.[90] propose a novel convolution and DDPM-based reconstruction model which leverages geometric priors to achieve fast generation of high-fidelity 3D textured meshes from a single image.
@@
Compared to generating textures from a single image, generating textures from multi-view images imposes higher demands on consistency between the images. Zhang et al.[91] propose a novel approach to achieve multi-view consistency. This method is based on a pre-trained diffusion model, such as DDIM, which decodes the predicted noise-free latent states into color space at each denoising step and blends them to form a texture image. An optimization process is then used to update the latent codes, ensuring that the images rendered from the blended texture image remain consistent across all views. Liu et al.[92] first fine-tunes a 2D diffusion model to predict a consistent set of images from six viewpoints, providing coherent views for 3D reconstruction. It then employs a 3D diffusion model under multi-view conditions to predict a textured mesh from the consistent multi-view images, using a coarse-to-fine approach. By leveraging multi-view images as supervision and applying lightweight optimization techniques, the method enhances the texture quality of the generated mesh.
@@
In addition to RGB color, lighting information is also an important component of texture. To generate shading results under specific lighting conditions, Zhang et al.[93] propose a new diffusion model that takes into account the given lighting environment during training. By applying the same ambient lighting during the material extraction process, DreamMat can produce high-quality PBR materials that are consistent with the given geometry and free from any baked shadow effects. On the contrary, Zeng et al.[94] introduce Paint3D, which aims to generate high-quality textures without embedded lighting information. By utilizing a pretrained 2D image diffusion model and a specially designed UV completion and upscaling diffusion model, Paint3D achieves the goal of generating high-quality, lighting-free 3D texture maps from text or image inputs.
To generate visually consistent textures, some works attempt to incorporate depth information into the generation process. Chen et al.[95] employ a pretrained depth-aware image diffusion model, which can generate high-resolution partial textures based on text prompts. Additionally, the authors progressively synthesize high-resolution local textures from different viewpoints to avoid the accumulation of inconsistencies and stretched artifacts between views. And Richardson et al.[96] use a pretrained depth-to-image diffusion model to render 3D models from different viewpoints, incorporating a dynamic trimap segmentation technique for generating seamless textures across various views. Moreover, unlike previous methods, Chen et al.[97] formulate the texture synthesis task as an optimization problem in RGB space. By employing a multi-resolution texture field and a cross-attention decoder, it effectively leverages the depth-to-image diffusion prior to generate high-quality, style-consistent textures for indoor scenes.
@@ IV-E Human Avatar Generation
Human avatar (head) generation is currently the most widely researched direction of human generation. Wang et al.[98] effectively generate detailed 3D digital avatars represented in NeRFs by unfolding 3D feature maps onto a 2D plane and applying 3D-aware convolutions which accounts for the projection relationship of 3D data onto the 2D plane. Jiang et al.[99] use diffusion models as guidance to generate 3D human avatars with parameterized shape and pose control based on text prompts. And Chen et al.[100] present a novel 3D deformation diffusion model capable of generating 3D mesh models based on input parameters such as identity and expression. By combining the 3D deformation model with a multi-view consistency diffusion approach, the model achieves the generation of controllable and realistic 3D avatars from a single image. Lei et al.[101] propose the DiffusionGAN3D framework, which combines a pretrained 3D generative model (such as EG3D) with a text-to-image diffusion model. The framework employs Relative Distance Loss to enhance diversity during the domain adaptation process and utilizes case-specific learnable triplanes to improve the network’s generative capabilities. Zhou et al.[102] introduce a new method called UltrAvatar, which uses Diffuse Color Extraction Model and Authenticity Guided Texture Diffusion Model to generate realistic and animatable 3D avatars from text prompts or a single image. To address several key issues including low texture quality, identity information loss, and inadequate occlusions handling, Li et al.[88] propose UV-IDM, which leverages the powerful texture generation capabilities of latent diffusion models (LDM) to generate realistic facial textures based on the Basel Face Model.
Compared to only generating an avatar, generating a human 3D model with a specific pose is more complex and challenging. Chen et al.[103] introduce the first diffusion model-based framework for 3D human generation called PrimDiffusion, which based on volumetric primitives, modeling the human body as multiple small volumes with radiance and motion information. By directly manipulating the denoising and diffusion processes on the volumetric primitives, the framework addresses the high computational cost of 3D representations and the complexity of 3D human topology. Subsequently, Kim et al.[104] present a 3D human generation pipeline called Chupa, aimed at addressing the challenge of generating realistic 3D human meshes with diverse identities, poses, and random details. They break the problem down into 2D normal map generation and normal map-based 3D reconstruction, improving computational efficiency and generating multi-view consistent 3D models. Svitov et al.[105] propose DINAR, which combines neural textures with the SMPL-X human model [153], achieving realistic quality while keeping the cartoon characters easy to manipulate and enabling fast inference. A latent diffusion model is used to recover the textures. Ho et al.[106] introduces a new 3D human reconstruction pipeline called SiTH, which addresses the challenge of reconstructing realistic and detail-rich 3D human models from a single-view image. The SiTH method breaks down the single-view reconstruction problem into two subproblems: generating the illusion of a back view image and mesh reconstruction. First, an image-conditioned diffusion model is used to predict the appearance of the back view of the input image. Then, using both the foreground and back view images as guidance, the method reconstructs a textured mesh of the entire body. Stathopoulos et al.[107] present a method called ScoreHMR, aimed at addressing the problem of predicting human model parameters from 2D image evidence. The ScoreHMR method mimics model fitting procedure, where a diffusion model is trained to capture the conditional distribution of human model parameters for a given input image. By using task-specific scores to guide its denoising process, ScoreHMR effectively resolves inverse problems across various applications without the need to retrain the diffusion model for each specific task. Huang et al.[108] present a new method called HumanNorm, which enhances the model’s 2D perception of 3D geometry by learning a normal-adapted diffusion model and a normal-aligned diffusion model. Additionally, it introduces a progressive geometry generation strategy and a multi-step Score Distillation Sampling (SDS) loss function to improve the performance of 3D human generation. Different from previous image-based methods, Kolotouros et al.[109] propose a method for generating realistic and animatable 3D human avatars solely from text descriptions. The authors leverage a large text-to-image diffusion model to generate an initial 3D human model, using NeRF to represent the 3D scene and optimize it to produce dynamic 3D human avatars with high-quality textures and learned instance-specific surface deformations.
@@
With the generated 3D human model, some studies go further to explore how to use DDPMs to generate a sequence of continuous actions. Gerogiannis et al.[110] propose a novel approach that integrates GNNs as denoising diffusion models, conducting the diffusion process directly in the mesh space to generate 3D facial expressions with a consistent noise sampling method, to ensure temporal coherence in animations. Xu et al.[111] introduce a new task of predicting human-object interactions (HOI) in 3D space, with the primary goal of generating long-term, diverse, and vivid 3D HOI sequences. The authors propose a framework called InterDiff, which combines interaction diffusion with physics-informed interaction correction to generate realistic and physically plausible future 3D sequences of human-object interactions. Since traditional motion diffusion models often neglect physical laws, leading to the generation of physically unrealistic motions, Yuan et al.[112] introduce a novel physics-guided human motion diffusion model called PhysDiff. The model incorporates physical constraints during the diffusion process by using a physics-based motion projection module which employs motion imitation from a physics simulator to project the denoised motions in the diffusion steps onto physically plausible movements.